Accelerator Update for December 3, 2007 - December 5, 2007
Monday December 3
The day shift began with Operations monitoring the Tevatron (TeV) store 5770 (81.48E30) and the Recycler stash (156.3E10), with the Antiproton Source (Pbar) stacking (35.7E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE experiments taking beam.
At 8:43 AM, the NuMI beam permit tripped off due to bad beam positions. It reset.
MI experts began aperture scans at MI-20 at 11:04 AM.
At 11:06 AM, Operations reported a Pbar emittance problem; a blowup had caused 2.8 mA of antiprotons to be lost. Operators investigated and found that experts working on the Pbar damper system had accidentally applied a test signal on chambers that affected the beam.
At 11:07 AM, two Booster RF stations (BRF2 & 4) tripped off. They both reset.
At 11:21 AM, the MI beam permit tripped off due to a kicker trip (KPS1B). Everything reset.
A Booster expert reported at 2:01 PM that BRF3 had been repaired and brought on line.
TeV experts began a 30-minute D0 alpha bump study at 3:28 PM.
Operations reported at 4:08 PM that the Recycler had stashed antiprotons three times during the day shift.
Operations terminated store 5770 around 5:06 PM.
Operations put the TeV through a wet squeeze at 6:09 PM.
Operations put the TeV into shot setup at 6:46 PM.
A TeV expert began a parasitic TeV study at 7:08 PM that consisted of taking special measurements of the beam during the shot setup.
At 7:20 PM, the Pbar beam permit dropped on a target station input for the first of two times in nine minutes. It reset both times.
At 9:15 PM, Operations reported that the TeV aborted during the scrape, one of the final steps in shot setup. There was no quench. Operators investigated.
Recycler experts began implementing the Recycler’s new damper system at 9:44 PM.
At 9:57 PM, Operations reported the loss of about 5E10 of antiprotons from the Recycler.
At 9:59 PM, BRF3 tripped off. It reset.
At 10:22 PM, Operations reported another loss of antiprotons from the Recycler stash. Operators contacted the expert working on the new damper system. This problem was caused by a discrepancy in how the application used system gain with the dampers on.
Operations prepared D0 for an access at 10:37 PM.
Recycler experts finished their damper work at 11:55 PM.
D0 came out of access at 11:55 PM.
Operations reported at 11:58 PM that the Recycler had stashed two times during the evening shift.
Tuesday December 4
The midnight shift began with Operations monitoring the stash (64.8E10), with Pbar stacking (72.1E10), and with NuMI and Sci /MiniBooNE taking beam.
Operations put the TeV through a dry squeeze at 12:25 AM.
Operations put the TeV into shot setup at 3:52 AM.
At 5:47 AM, the TeV suffered a quench at sector BA and BB. Operators investigated.
Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.
A cryo system expert reported at 8:05 AM that he was swapping a control card for the TeV sector A3 thermometry crate.
A Central Helium Liquefier (CHL) expert replaced a bad board in the TeV sector A3 thermometry crate at 8:31 AM.
At 8:35 AM, a MI kicker (I:KPS1B) tripped off for the first of two times in seven minutes. It reset. Later, an expert investigated and made adjustments.
At 10:02 AM, MI tripped off due to a ground fault indication for the area in between MI-20 and MI-40.
Due to the necessary access into MI, Recycler experts began a destructive study with the stash at 11:48 AM.
At 12:40 PM, Operations reported that the stash was gone and that they were preparing MI for access to search for the ground fault.
Experts in MI reported at 1:19 PM that they found a cable lying on the bus. They removed the cable and came out of access.
Operations established beam to Sci/MiniBooNE at 3:38 PM.
Operations established beam to NuMI and Pbar at 3:54 PM.
Operations reported at 4:27 PM that the Recycler had stashed once during the day shift.
Operations began loading the TeV with protons around 4:53 PM for tuning.
At 6:13 PM, a MI RF station (MIRF4) tripped off for the first of two times in 15 minutes. It reset.
Operations prepared D0 for access at 9:44 PM.
Vacuum experts and others accessed Linac at 10:45 PM. They reported that the leak was somewhere on the line leading to the Linac Momentum dump.
Operations put the TeV through a dry squeeze at 10:56 PM.
Operations reported at 11:55 PM that the Recycler had stashed two times during the evening shift.
Wednesday December 5
The midnight shift began with Linac in access, with Operations monitoring a stash (86.8E10) and stack (29.0E10), and with NuMI and Sci/MiniBooNE waiting on the Linac vacuum problems to be resolved.
Vacuum technicians reported at 1:20 AM that they found the vacuum leak on a seal on the first wire scanner just downstream of a Lambertson (B:LAM). The techs had the leak repaired and were out of Linac by 2:29 AM.
Linac resumed sending beam to Pbar and NuMI at 4:46 AM.
Operations reestablished beam to Sci/MiniBooNE at 5:36 AM.
At 6:25 AM, Operations reported that the Linac vacuum was having problems recovering.
Operations put the TeV through a wet squeeze at 7:16 AM.
At 7:21 AM, the Linac suffered a vacuum burst. Operations kept the Linac beam off, with the exception of sending protons to the TeV, so that the vacuum could recover.
Operations put the TeV into shot setup at 7:50 AM.
Operations reported at 8:15 AM that the Recycler had stashed two times during the midnight shift.
The Plans for Wednesday
The plans for today are to establish a store and to repair the Linac vacuum. During this repair, the Run Coordinator will allow Booster to access their tunnel and replace 2 or 3 weak RF power amplifiers. Other accesses may be allowed depending on the jobs.
Comments and Suggestions
|last modified 12/05/2007 email Fermilab|
|Security, Privacy, Legal|